OpenStudy (eric_d):

A lift is moving down with acceleration of 3ms^-2.A ball is released 1.7m above the floor.Assuming that g=9.8ms^-2,how long will the ball take to hit the floor

OpenStudy (eric_d):

@Abhisar

OpenStudy (abhisar):

Do u know the answer, i mean do u have an answer key ?

OpenStudy (eric_d):

Maybe,I need to use s=ut+1/2at^2

OpenStudy (abhisar):

yeah i know that, just wanted to confirm my answers. Use the same equation and take a=9.8

OpenStudy (eric_d):

The given answer is 0.71s

OpenStudy (abhisar):

is ur question complete ?

OpenStudy (eric_d):

yes

OpenStudy (eric_d):

Something is incomplete, when substituting the value Nt sure, whether this is correct I take 9.8-3=a I can get the correct answer

OpenStudy (abhisar):

a will be 9.8+3

OpenStudy (eric_d):

I tried that earlier I'm getting 0.27 as my final answer

OpenStudy (abhisar):

oh yes i got it !

OpenStudy (abhisar):

You will have to use the concept of relative motion.

OpenStudy (abhisar):

Since both the lift and ball are in motion that too with different accelerations. You will have to apply relative motion concept. Acceleration of ball = 9.8 (since it is not in contact with lift), acceleration of lift = 3. So acceleration of ball relative to lift will be 9.8-3= 6.8

OpenStudy (abhisar):

now take a=6.8 and use the second equation to find the value of t

OpenStudy (eric_d):

Okay Understood t^2=0.5 t=0.71 Thanks @Abhisar

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)